OpenStudy (anonymous):

solve the system using elimination x=-4y-4; 5x=4y=12

OpenStudy (mathstudent55):

x = -4y - 4 5x = 4y = 12 There are 2 equal signs in the second equation. Can you fix it?

OpenStudy (anonymous):

x=-4y-4 5x=4y=12

OpenStudy (anonymous):

5x+4y=12

OpenStudy (mathstudent55):

That's better. x = -4y - 4 5x + 4y = 12 Let's add 4y to both sides of the first equation, and just copy the second equation below it: x + 4y = -4 5x + 4y = 12 Ok, so far?

OpenStudy (anonymous):

i dont understand